# GLOVER: Generalizable Open-Vocabulary Affordance Reasoning for Task-Oriented Grasping

### v1

@@ Abstract
Inferring affordable (i.e., graspable) parts of arbitrary objects based on human specifications is essential for robots advancing toward open-vocabulary manipulation. Current grasp planners, however, are hindered by limited vision-language comprehension and time-consuming 3D radiance modeling, restricting real-time, open-vocabulary interactions with objects. To address these limitations, we propose GLOVER, a unified Generalizable Open-Vocabulary Affordance Reasoning framework, which fine-tunes the Large Language Models (LLMs) to predict visual affordance of graspable object parts within RGB feature space. We compile a dataset of over 10,000 images from human-object interactions, annotated with unified visual and linguistic affordance labels, to enable multi-modal fine-tuning. GLOVER inherits world knowledge and common-sense reasoning from LLMs, facilitating more fine-grained object understanding and sophisticated tool-use reasoning. To enable effective real-world deployment, we present Affordance-Aware Grasping Estimation (AGE), a non-parametric grasp planner that aligns the gripper pose with a superquadric surface derived from affordance data. In evaluations across 30 real-world scenes, GLOVER achieves success rates of 86.0% in part identification and 76.3% in grasping, with speeds approximately 330 times faster in affordance reasoning and 40 times faster in grasping pose estimation than the previous state-of-the-art.
Figure 1: Left: Compared with LERF-TOGO [65], GLOVER eliminates the need for capturing multi-view images for time-consuming 3D rendering and explicit instruction of the grasping part using LLMs. GLOVER is capable of providing more accurate fine-grained affordance predictions. Right: We demonstrate the efficacy of GLOVER in handling complex scenarios involving compositional object understanding and task-aware tool using under targeted human instructions, including: (a) attributes and relations like color, size, material, (b) tool function reasoning according to action, (c) common tool using in complex scenes.
@@ 1 Introduction
Human beings have an inherent ability to manipulate objects by understanding natural language instructions, such as distinguishing between different object types, identifying objects’ locations, and determining which part to grasp based on the desired task. Motivated by this, research in robotic grasping has evolved from focusing on closed-set objects [68, 16, 49] to open-vocabulary methods [27, 65, 77, 61, 23, 40, 55, 21, 22]. However, these open-vocabulary grasping methods suffer from a lack of complex reasoning capabilities regarding object properties. They face challenges in locating objects with subtle linguistic distinctions (e.g., distinguishing between a red apple and a green apple as shown in Fig. 1) and determining task-specific graspable parts (e.g., pick up the pot versus open the pot as shown in Fig. 1). For example, LERF-TOGO [65] relies on offline large language models (LLMs) to infer the graspable part based on the task (hammer handle), as we show in the left of Fig. 1. Despite this, it still fails to accurately identify grasp locations due to limited vision-language comprehension. Additionally, many previous approaches [25, 65, 77, 22, 61] depend heavily on 3D radiance modeling [50, 24] to construct 3D feature fields, which is both time-consuming and restricts their generalizability to new scenes.
To tackle the aforementioned challenges, we present a unified Generalizable Open-Vocabulary Affordance Reasoning (GLOVER) framework for open-vocabulary robotic grasping. This proposed framework aims to address three questions:
@@
1. How to represent the graspable part? Large Language Models (LLMs) demonstrate strong open-vocabulary reasoning capabilities and it is promising to fine-tune LLMs to output visual affordance masks. To achieve this, we define the graspable region as a global visual affordance mask, inspired by the visual affordance inferring [20, 47, 17, 60] and reasoning segmentation task [31, 41, 64]. Unlike a binary mask, this affordance mask encodes a continuous probability map, representing the likelihood of grasping at various locations. We adopt the affordance mask representation for two reasons: (1) graspable parts are better represented as regions rather than single points, and (2) predicting global masks based on language input aligns more naturally with model decoding [39, 26, 31, 19]. To support this approach, we compile a dataset of over 10,000 human-object interaction images, unifying the linguistic annotations and visual affordance annotation using a Gaussian distribution centered on graspable points.
@@
2. How to empower the model with complex reasoning and world knowledge to handle linguistic instructions? With the collected dataset, we fine-tune the model to generate visual affordance masks. Specifically, the fine-tuning process takes images and language instructions as input, leveraging LLaVA-7B [39] to perform multi-modal encoding following [31]. The encoded affordance token, which aggregates both visual and linguistic features, is fed into an affordance decoder to output the desired affordance mask. This mask is then projected into 3D space as stereo affordance for downstream tasks. GLOVER offers three main advantages: (1) It can leverage extensive 2D affordance annotations from real-world tasks, overcoming the scarcity of 3D affordance labels. (2) Fine-tuning allows GLOVER to inherit world knowledge and common-sense reasoning from the base LLM, enabling compositional object understanding and task-aware tool using as shown in Fig. 1. (3) GLOVER performs real-time, open-vocabulary affordance reasoning without scene rendering, and generalizes to novel scenes effectively.
@@
3. How to deploy the model efficiently in real robots? We introduce an Affordance-Aware Grasping Estimation (AGE) module that estimates gripper poses based on the geometry of affordance regions for real-world deployment. AGE is a non-parametric method that outperforms learning-based grasping planners [11, 12] in both performance and efficiency (40 times faster), without reliance on additional training data. Inspired by [42, 71], AGE samples grasping poses within the affordance space, aligning the gripper with the superquadric surface derived from the stereo affordance geometry. This approach eliminates the need for conventional grasping planners, enabling direct, affordance-aware pose estimation.
Our GLOVER module achieves state-of-the-art performance in the affordance benchmark, surpassing previous methods by a significant margin. To validate real-world applicability, we deploy GLOVER in a physical robot system, testing its open-vocabulary grasping capabilities across 30 challenging object scenes. GLOVER demonstrates an average improvement of 31.3% in affordance reasoning success rate and 33.6% in grasping success rate compared to LERF-TOGO [65]. Additionally, it achieves approximately $\times$330 faster affordance reasoning and $\times$40 faster grasping pose estimation.
@@ 2.1 Open-Vocabulary Representation for Manipulation
Many recent studies have worked on language-guided robotic tasks like navigation [2, 29, 28] and manipulation [9, 40, 65, 78]. The intervention of language plays a positive role in policy learning [16, 49, 22], value functions estimation [1, 35] and visual perception [61, 77]. Among them, the open-vocabulary manipulation is one of the most significant research topics. Several recent works focus on integrating 2D foundation models with 3D feature fields to achieve a 3D semantic-aware representation for open-vocabulary tasks [61, 65, 67, 77]. These methods distill features from 2D foundation models like CLIP [63], DINO [4] as training objective for NeRF [50] or GaussianSplatting [24] to reconstruct 3D feature fields. However, methods based on the 3D distilled feature fields heavily rely on time-consuming radiance modeling and require re-rendering when changing scenes, which limits the generalizability to some extent. In this work, we directly finetune the 2D foundation models to generate the related affordable areas in a generalizable and open-vocabulary manner, which is robust to scenario changes.
@@ 2.2 Task-Oriented Grasping
Task-oriented grasping refers to grasping different parts of objects based on the tasks. Previous research solves the task via detecting related object parts [53, 7, 45, 21], modeling 3D point clouds for affordance grounding [70, 14, 38], or transferring grasps to new instances based on category [14, 23]. Recent works [21, 9, 70, 45, 65, 23, 34] leverage vision-language models to reason the object parts for grasping. LERF-TOGO [65] derives a rough 3D object mask using DINO [4] features to expand a relevant area locally. Subsequently, a LERF [25] query is conditioned on this mask to separate sub-parts of the object. Robo-ABC [23] reasons the objects’ grasping point by using CLIP [63] to retrieve objects that share semantic similarity from the affordance memory. ShapeGrasp [34] infers the contact points by prompting the large language models via Chain of Thought [73]. Almost all the above methods rely on labeled 3D part-affordance datasets or additional pre-trained grasp networks like GraspNet [11], AnyGrasp [12] to infer grasping pose. However, GLOVER can leverage extensive 2D affordance data, well-trained 2D foundation models, and does not require an additional grasp planner network to estimate poses, which is more efficient.
@@ 2.3 Visual Affordance Reasoning
Previous research infers the affordance from human-object interactions [18, 20, 47], scene understanding [65, 44, 61] and 3D point cloud grounding [15, 54, 74, 52]. Recently, the foundation models such as LLMs and vision-language models (VLMs), have been integrated to perform affordance reasoning [17, 60, 23, 70, 45, 75]. AffCorrs [17] tackle one-shot visual affordance transfer by querying the object parts to find semantically corresponding ones via pre-trained DINO-ViT [4]. AffordanceLLM [60] train the LLaVA [39] on affordance dataset AGD20K [46], leveraging the world knowledge of the foundation model. Both Robo-ABC [23] and RAM [30] adopt the retrieve-and-transfer framework for zero-shot affordance reasoning. They construct the affordance memory from 2D images and retrieve the similar demonstration from the affordance memory with the help of CLIP [63] to reason affordance in the unseen domain. Our GLOVER does not require the creation of affordance memory, instead, it reasons the affordance leveraging LLM’s world knowledge in an end-to-end manner.
Figure 2: An overview of our method (Our contributions are highlighted with colored numbers). 1. We annotate the categories with the VLM and unify the affordance representation. 2. We fine-tune the affordance decoder to decode the affordance token [AFF], which encodes multi-modal information from multi-modal LLM. The fine-tuned GLOVER infers visual affordance in an open-vocabulary manner. 3. The affordance-aware grasping estimation module estimates gripper pose based on stereo affordance representation. See Fig. 3.
@@ 3.1 VL-Affordance Dataset Construction
Image Data collection. To enable affordance reasoning with the world knowledge of LLMs, we leverage the abundant resources of 2D images with visual affordance annotations. We select images from two common affordance datasets, AGD20K [46] and 3DOI [58]. AGD20K is a semi-supervised affordance dataset that includes 23,816 images with 50 categories of objects and 36 categories of affordance, sourced mainly from COCO [37], HICO [5] and free-license websites. For our purposes, we only use the testset split of AGD20K with well-annotated visual affordance points and corresponding text labels.
@@ 4 Affordance-Aware Grasping Estimation
With the inferred visual affordance, the next question is determining how the agent can interact with objects based on these affordances to effectively manipulate them. This question is challenging as the projected stereo affordances (see 4.1) from the visual affordances have complicated and irregular geometric surfaces, making it difficult to identify a global optimal point to grasp. To address this, we propose an affordance-aware grasping estimation (AGE) method. We model these irregular geometric surfaces using superquadric recovery techniques [42, 71, 69, 32, 6, 56, 57] and estimate the gripper pose by aligning it with the affordance-derived superquadric. We describe the details below (and illustrated in Fig. 3).
Figure 3: Illustration of Affordance-Aware Grasping Estimation (AGE). (a) Voxel down-sampled point clouds. (b) Filter the noise with DBSCAN [66] clustering. (c) Recover superquadric $\mathcal{A}$ from filtered stereo affordance. (d) Denote the gripper similarly as an ellipsoid surface $\mathcal{G}$. (e) Estimate the grasp pose by aligning the $\mathcal{A}$ and $\mathcal{G}$.
@@ 4.1 Stereo Affordance Preprocessing
We first map the deduced visual affordance onto stereo space using the RGB-D camera’s intrinsic parameters, before pre-processing the stereo affordance regions to reduce noise. Since the model may infer multiple clusters in the stereo affordance space, we filter out those with low affordance weights. Specifically, we first down-sample the point cloud with voxel down-sampling. Then, we apply DBSCAN [66], a density-based clustering algorithm, to divide the affordances into distinct clusters. We calculate the mean affordance weight for each cluster and discard the 3D point clouds of clusters with low mean weights.
@@ 4.2 Superquadric Recovery from Stereo Affordance
Superquadrics represent a class of geometric shapes that can model diverse forms. To determine the grasping pose from complex affordance regions, we reconstruct superquadrics from high-weight geometric structures within the affordance point clouds, representing each superquadric as $\mathcal{A}$. This process centers on identifying an optimal set of parameters $\boldsymbol{\lambda}\in\mathbb{R}^{11}$ to maximize alignment between $N$ stereo affordance points $\boldsymbol{a}_{i}=[x_{i},y_{i},z_{i}]$ ($i=1,\ldots,N$) and the superquadric surface.
@@
The affordance weight $\mathcal{W}_{i}$ ensures the inclination towards affordance points with high probability when estimating surface geometries of superquadrics. The inside-outside function $\left(F\left(\boldsymbol{a}_{i},\boldsymbol{\lambda_{\mathcal{A}}}\right)-1\right)^{2}$ aims to minimize the radial Euclidean distance, while the term $\lambda_{V}$ is the superquadric volume coefficient, deprecating the expansion of the superquadric volume. To further control the range of the recovered superquadric, we construct a penalty term based on the estimated volume of superquadrics, termed as $\beta V(\boldsymbol{\lambda_{\mathcal{A}}})$.

### v2

@@ Abstract
Inferring affordable (i.e., graspable) parts of arbitrary objects based on human specifications is essential for robots advancing toward open-vocabulary manipulation. Current grasp planners, however, are hindered by limited vision-language comprehension and time-consuming 3D radiance modeling, restricting real-time, open-vocabulary interactions with objects. To address these limitations, we propose GLOVER, a unified Generalizable Open-Vocabulary Affordance Reasoning framework, which fine-tunes the Large Language Models (LLMs) to predict the visual affordance of graspable object parts within RGB feature space. We compile a dataset of over 10,000 images from human-object interactions, annotated with unified visual and linguistic affordance labels, to enable multi-modal fine-tuning. GLOVER inherits world knowledge and common-sense reasoning from LLMs, facilitating more fine-grained object understanding and sophisticated tool-use reasoning. To enable effective real-world deployment, we present Affordance-Aware Grasping Estimation (AGE), a non-parametric grasp planner that aligns the gripper pose with a superquadric surface derived from affordance data. In evaluations across 30 table-top real-world scenes, GLOVER achieves success rates of 86.0% in part identification and 76.3% in grasping, with speeds approximately 29 times faster in affordance reasoning and 40 times faster in grasping pose estimation than the previous state-of-the-art. We also validate the generalization across embodiments, showing effectiveness in humanoid robots with dexterous hands.
Figure 1: Top: Compared with previous methods, GLOVER eliminates the need for capturing multi-view images for 3D rendering, explicit instruction of the grasping part, and building extra off-line affordance memory. GLOVER is capable of providing more accurate fine-grained affordance predictions. Mid: We demonstrate the efficacy of GLOVER in handling complex scenarios involving compositional object understanding and task-aware tool using under targeted human instructions, including: (a) attributes and relations like color, size, material, (b) tool function reasoning according to action, (c) common tool using in complex scenes. Bottom: We validate the effectiveness across embodiments (in a humanoid robot with dexterous hands), the robot grasps the oven handle based on inferred affordance.
@@ 1 Introduction
Human beings have an inherent ability to manipulate objects by understanding natural language instructions, such as distinguishing between different object types, identifying objects’ locations, and determining which part to grasp based on the desired task. Motivated by this, research in robotic grasping has evolved from focusing on closed-set objects [71, 16, 52] to open-vocabulary methods [29, 68, 80, 64, 25, 43, 58, 22, 23]. The previous methods mainly consisted of two categories, based on 3D radiance modeling [80, 68, 23, 64] and object retrieving [32, 25] (the top row of Fig. 1). Methods based on 3D radiance modeling necessitate the acquisition of images from multiple cameras and the rendering of each scene, which is time-consuming. On the other hand, approaches based on object retrieving require the establishment of an additional offline object memory, which is labor-intensive. Also, these methods suffer from a lack of complex reasoning capabilities regarding object properties. They face challenges in locating objects with subtle linguistic distinctions (e.g., distinguishing between a red apple and a green apple as shown in Fig. 1) and determining task-specific graspable parts (e.g., pick up the pot versus open the pot as shown in Fig. 1).
@@
In this work, we present a unified Generalizable Open-Vocabulary Affordance Reasoning (GLOVER) framework for open-vocabulary robotic grasping in an end-to-end manner. We aim to leverage the open-vocabulary reasoning capabilities of Large Language Models (LLMs) and fine-tune LLMs to output visual affordance masks. To achieve this, we define the graspable region as a global visual affordance mask, inspired by the visual affordance inferring [21, 50, 18, 63] and reasoning segmentation task [33, 44, 67]. Unlike a binary mask, this affordance mask encodes a continuous probability map, representing the likelihood of grasping at various locations. We adopt the affordance mask representation for two reasons: (1) graspable parts are better represented as regions rather than single points, and (2) predicting global masks based on language input aligns more naturally with model decoding [42, 28, 33, 20]. To support this approach, we collect over 10,000 human-object interaction images, using Vision-Language Models (VLMs) to annotate the language labels and unified Gaussian distribution to annotate visual affordance masks.
@@
With the collected dataset, we take the pairs of image and language instruction as input, leveraging LLaVA-7B [42] to perform multi-modal encoding following [33]. The encoded affordance token, which aggregates both visual and linguistic features, is fed into an affordance decoder to output the desired affordance mask. This mask is then projected into 3D space as stereo affordance for downstream tasks. The fine-tuning pipeline of GLOVER offers two key benefits: (1) It can leverage extensive 2D human-object interaction images from diverse kinds of datasets [17, 8, 61, 49], overcoming the scarcity of 3D affordance data. (2) The fine-tuning allows GLOVER to inherit world knowledge and common-sense reasoning from the base LLM, enabling compositional object understanding and task-aware tool using in an open-vocabulary manner, as shown in Fig. 1.
@@
To enable real-world grasping, we introduce an Affordance-Aware Grasping Estimation (AGE) module that estimates gripper poses based on the geometry of affordance regions. AGE is a non-parametric method that outperforms learning-based grasping planners [11, 12] in both performance and efficiency (40 times faster), without reliance on additional training data. Inspired by [45, 74], AGE samples grasping poses within the affordance space, determining the target pose by aligning the gripper with the superquadric surface derived from the stereo affordance geometry The alignment is optimized via nonlinear constrained optimization. This approach eliminates the usage of standalone grasping pose models [32, 25], enabling direct and affordance-aware pose estimation in an end-to-end manner.
To summarize, our contributions are as follows: (i) We present GLOVER, a unified end-to-end perception framework designed for open-vocabulary robotic grasping. To enhance its capabilities, we have curated and annotated a dataset of over 10,000 multi-modal affordance images for fine-tuning, enabling GLOVER to leverage world knowledge and common-sense reasoning inherited from large language models (LLMs). (ii) We propose an AGE module to perform non-parametric grasping pose estimation, which demonstrates a $\times$40 speed improvement over previous approaches. (iii) Our GLOVER module achieves state-of-the-art performance in the affordance benchmark, surpassing previous methods by a significant margin. We test the open-vocabulary grasping capability across 30 challenging table-top real-world scenes, demonstrating an average improvement of 20.0% in affordance reasoning success rate and 17.3% in grasping success rate compared to the previous state-of-the-art. We also demonstrate the generalization across diverse scenes and embodiments, showing effectiveness in 4 tasks with humanoid robots and dexterous hands.
@@ 2.1 Open-Vocabulary Representation for Manipulation
Many recent studies have worked on language-guided robotic tasks like navigation [2, 31, 30] and manipulation [9, 43, 68]. The intervention of language plays a positive role in policy learning [16, 52, 23], value functions estimation [1, 38] and visual perception [64, 80]. Among them, the open-vocabulary manipulation is one of the most significant research topics. Several recent works focus on integrating 2D foundation models with 3D feature fields to achieve a 3D semantic-aware representation for open-vocabulary tasks [64, 68, 70, 80]. These methods distill features from 2D foundation models like CLIP [66], DINO [4] as training objective for NeRF [53] or GaussianSplatting [26] to reconstruct 3D feature fields. In this work, we directly finetune the 2D foundation models to generate the related affordable areas in a generalizable and open-vocabulary manner, which is robust to scenario changes.
@@ 2.2 Task-Oriented Grasping
Task-oriented grasping refers to grasping different parts of objects based on the tasks. Previous research solves the task via detecting related object parts [56, 7, 48, 22, 36], modeling 3D point clouds for affordance grounding [73, 14, 41], or transferring grasps to new instances based on category [14, 25]. Recent works [22, 9, 73, 48, 68, 25, 37] leverage vision-language models to reason the object parts for grasping. LERF-TOGO [68] derives a rough 3D object mask using DINO [4] features to expand a relevant area locally. Subsequently, a LERF [27] query is conditioned on this mask to separate sub-parts of the object. Robo-ABC [25] reasons the objects’ grasping point by using CLIP [66] to retrieve objects that share semantic similarity from the affordance memory. ShapeGrasp [37] infers the contact points by prompting the large language models via Chain of Thought [76]. Almost all the above methods rely on labeled 3D part-affordance datasets or additional pre-trained grasp networks like GraspNet [11], AnyGrasp [12] to infer grasping pose. However, GLOVER can leverage extensive 2D affordance data, well-trained 2D foundation models, and does not require an additional grasp planner network to estimate poses, which is more efficient.
@@ 2.3 Visual Affordance Reasoning
Previous research infers the affordance from human-object interactions [19, 21, 50], scene understanding [68, 47, 64] and 3D point cloud grounding [15, 57, 77, 55]. Recently, the foundation models such as LLMs and VLMs, have been integrated to perform affordance reasoning [18, 63, 25, 73, 48, 78]. AffCorrs [18] tackle one-shot visual affordance transfer by querying the object parts to find semantically corresponding ones via pre-trained DINO-ViT [4]. AffordanceLLM [63] train the LLaVA [42] on affordance dataset AGD20K [49], leveraging the world knowledge of the foundation model. Both Robo-ABC [25] and RAM [32] adopt the retrieve-and-transfer framework for zero-shot affordance reasoning. They construct the affordance memory from 2D images and retrieve the similar demonstration from the affordance memory with the help of CLIP [66] to reason affordance in the unseen domain. Our GLOVER does not require the creation of affordance memory, instead, it reasons the affordance leveraging LLM’s world knowledge in an end-to-end manner.
Figure 2: An overview of our method (Our contributions are highlighted with colored numbers). 1. We annotate the categories with the VLM and unify the affordance representation. 2. We fine-tune the affordance decoder to decode the affordance token [AFF], which encodes multi-modal information from multi-modal LLM. The fine-tuned GLOVER infers visual affordance in an open-vocabulary manner. 3. The affordance-aware grasping estimation module (AGE), including (a) Voxel down-sampled point clouds. (b) Filter the noise with DBSCAN [69] clustering. (c) Recover superquadric $\mathcal{A}$ from filtered stereo affordance. (d) Denote the gripper similarly as an ellipsoid surface $\mathcal{G}$. (e) Estimate the grasp pose by aligning the $\mathcal{A}$ and $\mathcal{G}$.
@@ 3.1 VL-Affordance Dataset Construction
Image Data collection. To enable affordance reasoning with the world knowledge of LLMs, we leverage the abundant resources of 2D images with visual affordance annotations. We select images from two common affordance datasets, AGD20K [49] and 3DOI [61]. AGD20K is a semi-supervised affordance dataset that includes 23,816 images with 50 categories of objects and 36 categories of affordance, sourced mainly from COCO [40], HICO [5] and free-license websites.
@@ 4 Affordance-Aware Grasping Estimation
With the inferred visual affordance, the next question is determining how the agent can interact with objects based on these affordances to effectively manipulate them. This question is challenging as the projected stereo affordances (see Sec.4.1) from the visual affordances have complicated and irregular geometric surfaces, making it difficult to identify a global optimal point to grasp. Superquadric recovery [45, 74, 72, 34, 6, 59, 60] is an effective method for estimating superquadric of irregular surfaces, which can be used to estimate the geometry of stereo affordance. Based on the estimated superquadric of affordance region, the grasping pose is calculated via nonlinear constrained optimization. We describe the details below (and illustrated in the third part of Fig. 2).
@@ 4.1 Stereo Affordance Preprocessing
We first map the deduced visual affordance onto stereo space using the RGB-D camera’s intrinsic parameters. Since the model may infer multiple clusters in the stereo affordance space, we filter out those with low affordance weights. Specifically, we first down-sample the point cloud with voxel down-sampling. Then, we apply DBSCAN [69], a density-based clustering algorithm, to divide the affordances into distinct clusters. We calculate the mean affordance weight for each cluster and discard the 3D point clouds of clusters with low mean weights.
@@ 4.2 Superquadric Recovery from Stereo Affordance
Superquadrics represent a class of geometric shapes that can model diverse forms. To determine the grasping pose from complex affordance regions, we reconstruct superquadrics from the geometric structures of the affordance point clouds, representing each superquadric as $\mathcal{A}$. This process centers on identifying an optimal set of parameters $\boldsymbol{\lambda}\in\mathbb{R}^{11}$ to maximize alignment between $N$ stereo affordance points $\boldsymbol{a}_{i}=[x_{i},y_{i},z_{i}]$ ($i=1,\ldots,N$) and the superquadric surface.
@@
The affordance weight $\mathcal{W}_{i}$ ensures the inclination towards affordance points with high probability when estimating surface geometries of superquadrics. The inside-outside function $\left(F\left(\boldsymbol{a}_{i},\boldsymbol{\lambda_{\mathcal{A}}}\right)-1\right)^{2}$ aims to minimize the radial Euclidean distance, while the term $\lambda_{V}$ is the superquadric volume coefficient, deprecating the expansion of the superquadric volume. To further control the range of the recovered superquadric, we construct a penalty term based on the estimated volume of superquadrics, termed as $\beta V(\boldsymbol{\lambda_{\mathcal{A}}})$ to ensure the robustness of the method for noisy point clouds.